OpenStudy (anonymous):

Word problem help? The length of a rectangle is 3 yd less than twice the width, and the area of the rectangle is 65 yd^2. Find the dimensions of the rectangle.

OpenStudy (anonymous):

Ok so what you have here is a system of equations\[l=2w-3\] and\[lw=65\] I suggest using substitution and putting in (2w-3) for l in the area equation

OpenStudy (anonymous):

I'm not sure I understand.

OpenStudy (anonymous):

Well you know that\[l=2w-3\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

So can't you put in 2w-3 for l in the second equation? and make it\[w(2w-3)=65?\]

OpenStudy (anonymous):

So solve for w?

OpenStudy (anonymous):

Yes you are going to have to solve this like a quadratic

OpenStudy (anonymous):

w= -5 ?

OpenStudy (anonymous):

Well can you have a negative value in a rectangle? Remember this is a quadratic so I suggest finding the positive zero

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (anonymous):

Well -5 was one of the zeroes of the quadratic, how did you get that?

OpenStudy (anonymous):

I simplified both sides of the equation, subtracted 65 from both sides, factored out the left side of the equation, and set the factors equal to 0

OpenStudy (anonymous):

Ok so what were your two factors?

OpenStudy (anonymous):

You're really close

OpenStudy (anonymous):

(2w−13)(w+5)=0

OpenStudy (anonymous):

Ok so you got -5 by setting w+5 equal to zero What would you get when you set 2w-13 equal to zero?

OpenStudy (anonymous):

13/2

OpenStudy (anonymous):

or 6.5

OpenStudy (anonymous):

YEP well done :) So now substitute that either into l=2w-3 or lw=65 and solve for l

OpenStudy (anonymous):

so width is 6.5 and length would be 10?

OpenStudy (anonymous):

Yep that would be correct Good job :)